# FeDaL: Federated Dataset Learning for Time Series Foundation Models

### v1

@@ Appendix B Implementation Details
FL Methods
FedAvg AISTATS’17 [53]
FedProx MLSys’20 [41]
FedPer arXiv’19 [3]
FedRep ICML’21 [16]
FFTS AAAI’25 [9]
@@
Deep Models and Traditional Machine Learning Methods
TimeMixer [61] ICLR’24 [61]
GPT4TS NeurIPS’23 [82]
PatchTST ICLR’23 [54]
TimesNet ICLR’23 [66]
DLinear AAAI’23 [73]
Fedformer ICML’22
Autoformer NeurIPS’21
Stationary NeurIPS’22 [47]
LightTS arXiv’22 [78]
Informer AAAI’21 [80]
Reformer ICLR’20 [36]
Time-LLM ICLR’24 [34]
N-HiTS AAAI’23 [6]
N-BEATS ICLR’20 [55]
ETSformer arXiv’22 [65]
MTCN ICLR’24 [51]
Pyraformer ICLR’22 [45]
Anomaly ICLR’22 [69]
DTW ICML’17 [17]
TS2Vec AAAI’22 [72]
T-Loss NeurIPS’19 [26]
TNC ICLR’21 [59]
TS-TCC IJCAI’21 [24]
@@
MCDNN CVPR’12 [15]
ResNet CVPR’16 [31]
t-LeNet ECML’16 [38]
TWIESN TR’06 [50]
@@
XGBoost KDD’16 [14]
Rocket DMKD’20 [21]
LSTM NC’97 [33]
LSTNet SIGIR’18 [37]
LSSL ICLR’22 [29]
Transformer NeurIPS’17 [60]
Flowformer ICML’22 [67]
LogTransformer NeurIPS’19 [39]
TCN NeurIPS’19 [26]
Foundation Model
Moment ICML’24 [28]
Moiraismall ICML’24 [64]
Morialbase ICML’24 [64]
Moriallarge ICML’24 [64]
TimesFM ICML’24 [19]
Chronossmall TMLR’24 [2]
Chronosbase TMLR’24 [2]
Chronoslarge TMLR’24 [2]
We compare our proposed FeDaL against 54 different baselines in our experiments to demonstrate its effectiveness and superiority in federated time series representation learning and downstream time series analysis tasks. In this section, we introduce some of these representative baselines from both federated learning and deep time series modeling. Full information of baselines can be found in Table˜12. The key federated learning baselines include:
  • FedAvg [53]: A decentralized approach that enables devices to collaboratively learn a shared model by aggregating locally-computed updates, allowing for the training of deep networks on private and large datasets while reducing communication costs.
  • FedProx [41]: A federated learning framework that generalizes and re-parameterizes FedAvg to tackle systems and statistical heterogeneity, providing convergence guarantees and demonstrating more robust and accurate convergence behavior.
  • FedPer [3]: A federated learning baseline that personalizes models by keeping client-specific layers local while training shared base layers collaboratively.
  • FedRep [16]: A federated learning baseline that decouples learning into global feature representation (shared) and client-specific heads (local) for better personalization.
  • FFTS [9] A FL-based decentralized framework for training time series foundation models from scratch, which incorporates a Mixture-of-Experts mechanism on each client and employs a heterogeneous knowledge alignment strategy to address time series heterogeneity.
@@
  • LogTransformer [39]: A modified Transformer architecture that addresses the locality-agnostics and memory bottleneck issues in time series forecasting by incorporating convolutional self-attention and log-sparse. attention, achieving improved forecasting accuracy with reduced memory cost.
@@
  • N-BEATS [55]: A deep neural architecture that achieves state-of-the-art performance in univariate time series point forecasting, using a stack of fully-connected layers with backward and forward residual links.
@@
  • Reformer [36]: This model improves Transformer by using locality-sensitive hashing for attention and reversible residual layers. It offers better memory efficiency and speed for lengthy sequences without sacrificing performance.
@@
  • Informer [80]: An optimized Transformer-based model for long-range time series prediction. It uses ProbSparse self-attention for efficiency, processes long inputs effectively, and employs a fast prediction decoder.
@@
  • LightTS [78]: A lightweight MLP structure. It utilizes two downsampling strategies—spaced and sequential sampling—on the MLP structure, capitalizing on the fact that downsampled time series generally maintain most of their original information.
@@
  • ETSformer [65]: A Transformer architecture that leverages the principle of exponential smoothing to improve traditional Transformers for time-series forecasting, offering better decomposition capability, interpretability, and long-term forecasting efficiency.
@@
  • Stationary (NS*) [47]: A framework that addresses the over-stationarization problem in time series forecasting by combining Series Stationarization and De-stationary Attention modules, which unify statistics for better predictability while recovering intrinsic non-stationary information.
@@
  • Autoformer [68]: A decomposition architecture that leverages leveraging an Auto-Correlation mechanism for long-term forecasting, which efficiently discovers dependencies and aggregates representations at the sub-series level.
@@
  • FEDformer [81]: A forecasting method that combines seasonal-trend decomposition with a frequency-enhanced Transformer, capturing both the global profile and detailed structures of time series.
@@
  • Pyraformer [45]: It features hierarchical pyramidal attention modules with binary trees to capture temporal dependencies across different ranges efficiently, both in time and memory complexity.
@@
  • AnomalyTransformer [69]: A approach for unsupervised time series anomaly detection that leverages the self-attention mechanism to compute association discrepancy, which captures the adjacent-concentration bias of anomalies.
@@
  • TimesNet [66]: A framework that transforms 1D time series into 2D tensors to model complex temporal variations, leveraging the multi-periodicity of time series to adaptively discover and extract intraperiod- and interperiod-variations.
@@
  • PatchTST [54]: This method divides the time series into patches at the sub-series level for input to the Transformer. Each channel holds a univariate time series, sharing the same embedding and Transformer weights across all series.
@@
  • DLinear [73]: DLinear integrates decomposition schemes from Autoformer and FEDformer with linear layers to model time series data tables. It effectively summarizes trend and seasonal components, enhancing performance on datasets rich in trends.
@@
  • N-HiTS [6]: A forecasting model that addresses the challenges of long-horizon forecasting by incorporating hierarchical interpolation and multi-rate data sampling techniques.
@@
  • GPT4TS [82]: This model is designed for time series analysis across various scenarios, achieved by fine-tuning a pre-trained language model, specifically GPT2, for the time series domain.
@@
  • Time-LLM [34]: A reprogramming framework that repurposes large language models (LLMs) for general time series forecasting by aligning time series data with natural language modalities through text prototypes and Prompt-as-Prefix (PaP) techniques.
@@
  • TimeMixer [61]: An advanced time series forecasting baseline that employs a multi-scale mixing architecture with Past Decomposition Mixing (PDM) and Future Multi-predictor Mixing (FMM) modules to effectively disentangle and integrate seasonal and trend patterns.
@@
OWA $=\frac{1}{2}\left[\frac{\text{SMAPE}}{\text{SMAPE}_{\textrm{Na\"{\i}ve2}}}+\frac{\text{MASE}}{\text{MASE}_{\textrm{Na\"{\i}ve2}}}\right],$
@@ Appendix C Full Results
This section presents the complete experimental results across full-shot and zero-shot long-term forecasting, short-term forecasting, imputation, anomaly detection, and classification. It also includes detailed plots and analysis of federated scaling behaviors during time series foundation model training.
C.1 Long-term Forecasting
The full long-term forecasting results are provided in Table˜13 and Table˜14. Table˜13 demonstrates that FeDaL-trained TSFM consistently outperforms state-of-the-art deep time series models specifically designed for long-term forecasting. Furthermore, Table˜14 shows that under zero-shot evaluation, our approach surpasses both advanced centralized pretrained TSFMs and federated pretraining baselines, highlighting its strong generalization across domains and tasks.

### v2

@@ Appendix B Implementation Details
Data Processing Pipeline
Given the varying lengths of time series in the pretraining dataset, some sequences are too short to support the full sequence length used during pretraining (3072 points). Moreover, despite the large scale of LOTSA, we aim to expose the model to as diverse a set of temporal patterns as possible. To address these challenges, we introduce an efficient dataset processing pipeline for client-side pretraining. This pipeline (1) filters out invalid or insufficiently long time series samples, (2) augments the original time series to enrich temporal diversity, and (3) flexibly handles multi-domain time series with varying numbers of channels or variables in a unified manner. The source code for the unified client-side data processing pipeline is provided below.
1class FederatedDataset_Representation(Dataset):
2 def __init__(self, args, series_data: dict,
3 mode=’train’,
4 ls=’normal’,
5 scale=True,
6 split_ratio=0.7):
7 self.initialize_parameters(args, mode, ls, scale)
8 self.process_data(series_data, split_ratio)
9
10 def initialize_parameters(self, args, mode, ls, scale):
11 self.args = args
12 self.mode = mode
13 self.ls = ls
14 self.scale = scale
15
16 self.subset_size = args.subset_size
17 self.window_stride = args.window_stride
18 self.seq_len = args.seq_len
19 self.pred_len = args.pred_len
20 self.total_len = self.seq_len + self.pred_len
21
22 def process_data(self, series_data: dict, split_ratio: float):
23 processed_series = []
24 skipped_series = 0
25
26 items_to_process = list(series_data.items())
27 for key, series in items_to_process:
28 required_len = self.seq_len
29 if len(series) < required_len:
30 skipped_series += 1
31 continue
32
33 if self.scale:
34 scaler = StandardScaler()
35 series = scaler.fit_transform(series)
36
37 n_features = series.shape[1]
38 for feature_idx in range(n_features):
39 feature_series = series[:, feature_idx]
40
41 windows = self.create_windows(
42 feature_series.reshape(-1, 1),
43 required_len,
44 self.window_stride
45 )
46 processed_series.extend(windows)
47
48 if len(processed_series) == 0:
49 raise ValueError("No valid sequences found after length filtering!")
50
51 self.data = np.array(processed_series)
52 split_idx = int(len(self.data) * split_ratio)
53
54 def __len__(self):
55 return len(self.data)
56
57 def __getitem__(self, index):
58 seq_x = self.data[index]
59 return torch.from_numpy(seq_x)
60
61 @staticmethod
62 def create_windows(series, window_size, stride=32):
63 windows = []
64 for i in range(0, len(series) - window_size + 1, stride):
65 windows.append(series[i:i + window_size])
66 return windows
Data Partitioning in Federated Time Series Representation Learning
Learning domain-invariant representations from unlabeled, heterogeneous time series is crucial for advancing TSFMs in federated settings. To evaluate FeDaL in this context, we adopt two large-scale cross-domain benchmarks: UTSD (Woo et al., 2024) (1B time points across 7 domains) with domain-mixed (DM) partitioning, and CTSD (Chen et al., 2025) (500M time points across 6 domains) with domain-independent (DI) partitioning. In the DM setting, each client receives data drawn from multiple domains, while in the DI setting, each client is restricted to a single domain. For better illustration, examples of both partitioning strategies are shown in Figure 9. To further assess FeDaL under varying degrees of heterogeneity, we construct two additional DM-based configurations on UTSD, denoted H1 and H2. These are generated using a customized data-loading pipeline that controls heterogeneity through an imbalance factor (parameter: imbalance_factor) (set to 0.1 for H1 and 0.3 for H2) in the static method _split_series. This allows evaluation of FeDaL’s robustness across environments with increasing domain skew.
1class FedRepresentationDatasetConfig:
2 DATASET_REGISTRY: Dict[str, str] = {
3 "UTSD": "UTSD"
4 }
5 SINGLE_DATASET_CONFIGS: Dict[str, Dict] = {
6 "UTSD_unbalance_H2":{
7 "dataset": "UTSD",
8 "max_clients": 10,
9 "split_method": "random",
10 "balanced": False,
11 "imbalance_factor": 0.3 # by default
12 },
13 "UTSD_unbalance_H1":{
14 "dataset": "UTSD",
15 "max_clients": 10,
16 "split_method": "random",
17 "balanced": False,
18 "imbalance_factor": 0.1 # by default
19 }
20 }
21
22 @classmethod
23 def get_max_clients(cls, dataset_name: str) -> int:
24 if dataset_name not in cls.SINGLE_DATASET_CONFIGS:
25 raise ValueError(f"Dataset {dataset_name} not found in configurations")
26 return cls.SINGLE_DATASET_CONFIGS[dataset_name]["max_clients"]
27
28 @classmethod
29 def get_total_series(cls, dataset_name: str) -> int:
30 if dataset_name not in cls.SINGLE_DATASET_CONFIGS:
31 raise ValueError(f"Dataset {dataset_name} not found in configurations")
32
33 dataset_key = cls.SINGLE_DATASET_CONFIGS[dataset_name]["dataset"]
34 file_path = os.path.join(’datasets’, ’crossdomain’, f’{cls.DATASET_REGISTRY[dataset_key]}.npz’)
35
36 with np.load(file_path) as data:
37 return len(data.keys())
38
39 @classmethod
40 def get_series_client_data(cls, dataset_name: str, client_id: int):
41 if dataset_name not in cls.SINGLE_DATASET_CONFIGS:
42 raise ValueError(f"Dataset {dataset_name} not configured for series-based splitting")
43
44 config = cls.SINGLE_DATASET_CONFIGS[dataset_name]
45 if client_id >= config["max_clients"]:
46 raise ValueError(f"Client ID {client_id} exceeds maximum clients {config[’max_clients’]}")
47
48 dataset_key = cls.SINGLE_DATASET_CONFIGS[dataset_name]["dataset"]
49 file_path = os.path.join(’datasets’, ’crossdomain’, f’{cls.DATASET_REGISTRY[dataset_key]}.npz’)
50 split_method = config["split_method"]
51
52 return cls._split_series(file_path, split_method, client_id, config["max_clients"], config["balanced"], config["imbalance_factor"])
53
54 @staticmethod
55 def _split_series(file_path: str, split_method: str, client_id: int, max_clients: int, balanced: bool = True, imbalance_factor: float = 0.3):
56 with np.load(file_path) as data:
57 keys = list(data.keys())
58 n_series = len(keys)
59
60 if split_method == "random":
61 np.random.seed(42)
62 shuffled_keys = np.random.permutation(keys)
63
64 if balanced:
65 series_per_client = n_series // max_clients
66 extra_series = n_series % max_clients
67 start_idx = client_id * series_per_client + min(client_id, extra_series)
68 end_idx = start_idx + series_per_client + (1 if client_id < extra_series else 0)
69
70 else:
71 alpha = np.ones(max_clients) * (1.0 / imbalance_factor)
72 proportions = np.random.dirichlet(alpha)
73 series_counts = (proportions * n_series).astype(int)
74 series_counts[-1] = n_series - np.sum(series_counts[:-1])
75 start_idx = np.sum(series_counts[:client_id])
76 end_idx = start_idx + series_counts[client_id]
77 client_series = {}
78 for key in shuffled_keys[start_idx:end_idx]:
79 client_series[key] = data[key]
80 else:
81 raise ValueError(f"Unknown split method: {split_method}")
82
83 return client_series
@@ 5 SINGLE_DATASET_CONFIGS: Dict[str, Dict] = {
FL Methods FedAvg AISTATS’17 (2017)
FedProx MLSys’20 (2020)
FedPer arXiv’19 (2019)
FedRep ICML’21 (2021)
FFTS AAAI’25 (2025)
@@
Deep Models and Traditional Machine Learning Methods TimeMixer (2024b) ICLR’24 (2024b)
GPT4TS NeurIPS’23 (2023)
PatchTST ICLR’23 (2022)
TimesNet ICLR’23 (2022a)
DLinear AAAI’23 (2023)
Fedformer ICML’22 (2022)
Autoformer NeurIPS’21 (2021)
Stationary NeurIPS’22 (2022)
LightTS arXiv’22 (2022)
Informer AAAI’21 (2021)
Reformer ICLR’20 (2020)
Time-LLM ICLR’24 (2023)
N-HiTS AAAI’23 (2022)
N-BEATS ICLR’20 (2020)
ETSformer arXiv’22 (2022)
MTCN ICLR’24 (2024)
Pyraformer ICLR’22 (2021)
Anomaly ICLR’22 (2022)
DTW ICML’17 (2017)
TS2Vec AAAI’22 (2022)
T-Loss NeurIPS’19 (2019)
TNC ICLR’21 (2021)
TS-TCC IJCAI’21 (2021)
@@
MCDNN CVPR’12 (2012)
ResNet CVPR’16 (2016)
t-LeNet ECML’16 (2016)
TWIESN TR’06 (2006)
@@
XGBoost KDD’16 (2016)
Rocket DMKD’20 (2020)
LSTM NC’97 (1997)
LSTNet SIGIR’18 (2018)
LSSL ICLR’22 (2021)
Transformer NeurIPS’17 (2017)
Flowformer ICML’22 (2022b)
LogTransformer NeurIPS’19 (2019)
TCN NeurIPS’19 (2019)
Foundation Model Moment ICML’24 (2024)
Moiraismall ICML’24 (2024)
Morialbase ICML’24 (2024)
Moriallarge ICML’24 (2024)
TimesFM ICML’24 (2024)
Chronossmall TMLR’24 (2024)
Chronosbase TMLR’24 (2024)
Chronoslarge TMLR’24 (2024)
We compare our proposed FeDaL against 54 different baselines in our experiments to demonstrate its effectiveness and superiority in federated time series representation learning and downstream time series analysis tasks. In this section, we introduce some of these representative baselines from both federated learning and deep time series modeling. Full information of baselines can be found in Table 15. The key federated learning baselines include:
  • FedAvg (McMahan et al., 2017): A decentralized approach that enables devices to collaboratively learn a shared model by aggregating locally-computed updates, allowing for the training of deep networks on private and large datasets while reducing communication costs.
  • FedProx (Li et al., 2020): A federated learning framework that generalizes and re-parameterizes FedAvg to tackle systems and statistical heterogeneity, providing convergence guarantees and demonstrating more robust and accurate convergence behavior.
  • FedPer (Arivazhagan et al., 2019): A federated learning baseline that personalizes models by keeping client-specific layers local while training shared base layers collaboratively.
  • FedRep (Collins et al., 2021): A federated learning baseline that decouples learning into global feature representation (shared) and client-specific heads (local) for better personalization.
  • FFTS (Chen et al., 2025) A FL-based decentralized framework for training time series foundation models from scratch, which incorporates a Mixture-of-Experts mechanism on each client and employs a heterogeneous knowledge alignment strategy to address time series heterogeneity.
@@
  • LogTransformer (Li et al., 2019): A modified Transformer architecture that addresses the locality-agnostics and memory bottleneck issues in time series forecasting by incorporating convolutional self-attention and log-sparse. attention, achieving improved forecasting accuracy with reduced memory cost.
@@
  • N-BEATS (Oreshkin et al., 2020): A deep neural architecture that achieves state-of-the-art performance in univariate time series point forecasting, using a stack of fully-connected layers with backward and forward residual links.
@@
  • Reformer (Kitaev et al., 2020): This model improves Transformer by using locality-sensitive hashing for attention and reversible residual layers. It offers better memory efficiency and speed for lengthy sequences without sacrificing performance.
@@
  • Informer (Zhou et al., 2021): An optimized Transformer-based model for long-range time series prediction. It uses ProbSparse self-attention for efficiency, processes long inputs effectively, and employs a fast prediction decoder.
@@
  • LightTS (Zhang et al., 2022): A lightweight MLP structure. It utilizes two downsampling strategies—spaced and sequential sampling—on the MLP structure, capitalizing on the fact that downsampled time series generally maintain most of their original information.
@@
  • ETSformer (Woo et al., 2022): A Transformer architecture that leverages the principle of exponential smoothing to improve traditional Transformers for time-series forecasting, offering better decomposition capability, interpretability, and long-term forecasting efficiency.
@@
  • Stationary (Liu et al., 2022): A framework that addresses the over-stationarization problem in time series forecasting by combining Series Stationarization and De-stationary Attention modules, which unify statistics for better predictability while recovering intrinsic non-stationary information.
@@
  • Autoformer (Wu et al., 2021): A decomposition architecture that leverages leveraging an Auto-Correlation mechanism for long-term forecasting, which efficiently discovers dependencies and aggregates representations at the sub-series level.
@@
  • FEDformer (Zhou et al., 2022): A forecasting method that combines seasonal-trend decomposition with a frequency-enhanced Transformer, capturing both the global profile and detailed structures of time series.
@@
  • Pyraformer (Liu et al., 2021): It features hierarchical pyramidal attention modules with binary trees to capture temporal dependencies across different ranges efficiently, both in time and memory complexity.
@@
  • AnomalyTransformer (Xu et al., 2022): A approach for unsupervised time series anomaly detection that leverages the self-attention mechanism to compute association discrepancy, which captures the adjacent-concentration bias of anomalies.
@@
  • TimesNet (Wu et al., 2022a): A framework that transforms 1D time series into 2D tensors to model complex temporal variations, leveraging the multi-periodicity of time series to adaptively discover and extract intraperiod- and interperiod-variations.
@@
  • PatchTST (Nie et al., 2022): This method divides the time series into patches at the sub-series level for input to the Transformer. Each channel holds a univariate time series, sharing the same embedding and Transformer weights across all series.
@@
  • DLinear (Zeng et al., 2023): DLinear integrates decomposition schemes from Autoformer and FEDformer with linear layers to model time series data tables. It effectively summarizes trend and seasonal components, enhancing performance on datasets rich in trends.
@@
  • N-HiTS (Challu et al., 2022): A forecasting model that addresses the challenges of long-horizon forecasting by incorporating hierarchical interpolation and multi-rate data sampling techniques.
@@
  • GPT4TS (Zhou et al., 2023): This model is designed for time series analysis across various scenarios, achieved by fine-tuning a pre-trained language model, specifically GPT2, for the time series domain.
@@
  • Time-LLM (Jin et al., 2023): A reprogramming framework that repurposes large language models (LLMs) for general time series forecasting by aligning time series data with natural language modalities through text prototypes and Prompt-as-Prefix (PaP) techniques.
@@
  • TimeMixer (Wang et al., 2024b): An advanced time series forecasting baseline that employs a multi-scale mixing architecture with Past Decomposition Mixing (PDM) and Future Multi-predictor Mixing (FMM) modules to effectively disentangle and integrate seasonal and trend patterns.
@@
OWA $=\frac{1}{2}\left[\frac{\text{SMAPE}}{\text{SMAPE}_{\textrm{Naïve2}}}+\frac{\text{MASE}}{\text{MASE}_{\textrm{Naïve2}}}\right],$
@@ Appendix C Full Results
This section presents the complete results across full-shot and zero-shot long-term forecasting, short-term forecasting, imputation, anomaly detection, and classification. It also includes detailed plots and analysis of federated scaling behaviors during general time series foundation model training.
C.1 Time Series Representation Learning
The full federated time series representation learning results are shown in Table 16, which shows our proposed FeDaL outperform advanced FL methods across various datasets and setups in unsupervised representation learning.
Table 16: Representation learning results under varying patch masking ratios. For UTSD, we simulate two levels of domain heterogeneity (H1 and H2) as detailed in the Appendix B. For CTSD, each dataset is treated as a domain-independent client. Bold indicates the best result, Underline the second best. $\dagger$} denotes evaluation using personalized models after server averaging; $\ddagger$} indicates client-side evaluation without aggregation.
Method
USTD Dataset (10 Clients, DM)
CTSD Dataset (18 Clients, DI)
20%
35%
50%
75%
90%
20% 35% 50% 75% 90%
H1 H2 H1 H2 H1 H2 H1 H2 H1 H2
Representation Learning Under Increased Heterogeneity
. We further investigated FeDaL’s performance under increasingly challenging, high-heterogeneity configurations, specifically employing imbalance ratios of $\{0.7,0.8,0.9\}$. The results are shown in Table 17, illustrate FeDaL‘s superior generalization and robust performance over competitive baselines. This confirms its ability to effectively handle severe non-IID conditions and maintain its leading edge.
Table 17: Federated representation learning results across different masking ratios and imbalance ratios (H3: 0.7, H4: 0.8, H5: 0.9).
Imbalance Ratio Method Masking Ratio
(H) 20% 35% 50% 75% 90%
H3 (0.7) FedAvg 0.427 0.505 0.571 0.623 0.912
FedProx 0.409 0.500 0.568 0.627 0.900
FedPer 0.399 0.612 0.544 0.582 0.877
FedRep 0.415 0.579 0.540 0.603 0.868
FFTS 0.339 0.423 0.525 0.599 0.831
FeDaL (Ours) 0.310 0.429 0.501 0.557 0.809
H4 (0.8) FedAvg 0.431 0.506 0.582 0.626 0.917
FedProx 0.412 0.509 0.577 0.630 0.926
FedPer 0.406 0.618 0.560 0.597 0.886
FedRep 0.418 0.592 0.555 0.606 0.871
FFTS 0.350 0.435 0.537 0.601 0.844
FeDaL (Ours) 0.322 0.430 0.511 0.560 0.824
H5 (0.9) FedAvg 0.451 0.523 0.598 0.637 0.924
FedProx 0.420 0.514 0.586 0.642 0.933
FedPer 0.416 0.620 0.577 0.603 0.878
FedRep 0.432 0.604 0.569 0.610 0.878
FFTS 0.362 0.441 0.550 0.612 0.846
FeDaL (Ours) 0.330 0.437 0.523 0.571 0.825
C.2 Long-term Forecasting
The full long-term forecasting results are provided in Table 18 and Table 19. Table 18 demonstrates that FeDaL-trained TSFM consistently outperforms state-of-the-art deep time series models specifically designed for long-term forecasting. Furthermore, Table 19 shows that under zero-shot evaluation, our approach surpasses both advanced centralized pretrained TSFMs and federated pretraining baselines, highlighting its strong generalization across domains and tasks.